# Stylized facts in Web3

### v2

@@ Appendix 1. Tail Behavior Analysis
$\hat{F}(x)\equiv\frac{1}{n}\sum^{n}_{i=1}\mathbf{1}_{X_{i}\leq x}$
@@
We flip the $x<0$ part to the positive x-axis, enabling simultaneous examination of both tails. Figures 18, 19, 20, and 21 display Equation 4 for various assets and timescales.
@@
Figures 19, 20, and 21 extend the analysis to all other tokens in Table 1 by presenting the one-hour returns CDF (Equation 4). The tails of the CDF are power law with a coefficient of $\alpha\sim 2.8$, similar to the value in the main text.
@@
Figure 19. Log-log plot of CDF tails using Equation (4) for L1 tokens’ one-hour normalized returns. A power law fit $\sim|x|^{-\alpha}$ with $\alpha=2.8$ is overlaid for comparison.
Figure 20. Log-log plot of CDF tails using Equation (4) for miscellaneous tokens’ one-hour normalized returns. A power law fit $\sim|x|^{-\alpha}$ with $\alpha=2.8$ is included for reference.
Figure 21. Log-log plot of CDF tails using Equation (4) for DeFi tokens’ one-hour normalized returns. A power law fit $\sim|x|^{-\alpha}$ with $\alpha=2.8$ is shown for comparison.
@@ Appendix 2. Autocorrelation details
$\frac{\frac{1}{\sqrt{N}}\displaystyle\sum_{(R_{t},R_{t+k})\in I_{k}}(R_{t}-\hat{\mu}1)(R{t+k}-\hat{\mu}_{2})}{\hat{\sigma}_{1}\hat{\sigma}_{2}}$ (5)
@@
Equation 5 calculates ACF for every hour in a day, resulting in "gap" periods with no pairs of hours for SPY (e.g., at a lag of 12 hours).
Figures 22, 23, 24, 25, 26, and 27 display autocorrelations of returns and absolute returns for tokens listed in Table 1, using all available historical data. These figures support the main text findings: return autocorrelations are mostly noise except for specific lags, while absolute return autocorrelations are large and decay slowly, indicating volatility clustering.

### v3

@@ Appendix 1. Tail Behavior Analysis
$\hat{F}(x)\equiv\frac{1}{n}\sum^{n}_{i=1}\mathbf{1}_{X_{i}\leq x}$ (7)
@@
We flip the $x<0$ part to the positive x-axis, enabling simultaneous examination of both tails. Figures 20, 21, 22, and 23 display Equation (8) for various assets and timescales.
@@
Figures 21, 22, and 23 extend the analysis to all other tokens in Table 1 by presenting the one-hour returns CDF (Equation 8). The tails of the CDF are power law with a coefficient of $\alpha\sim 2.63$, similar to the value in the main text.
@@
Figure 21. Log-log plot of CDF tails using Equation (8) for L1 tokens’ one-hour normalized returns. A power law fit $\sim|x|^{-\alpha}$ with $\alpha=2.63$ is overlaid for comparison.
Figure 22. Log-log plot of CDF tails using Equation (8) for miscellaneous tokens’ one-hour normalized returns. A power law fit $\sim|x|^{-\alpha}$ with $\alpha=2.63$ is included for reference.
Figure 23. Log-log plot of CDF tails using Equation (8) for DeFi tokens’ one-hour normalized returns. A power law fit $\sim|x|^{-\alpha}$ with $\alpha=2.63$ is shown for comparison.
@@ Appendix 2. Autocorrelation details
$\frac{\frac{1}{\sqrt{N}}\displaystyle\sum_{(R_{t},R_{t+k})\in I_{k}}(R_{t}-\hat{\mu}_{1})(R_{t+k}-\hat{\mu}_{2})}{\hat{\sigma}_{1}\hat{\sigma}_{2}}$ (9)
@@
Equation (9) calculates ACF for every hour in a day, resulting in "gap" periods with no pairs of hours for SPY (e.g., at a lag of 12 hours).
Figures 24, 25, 26, 27, 28, and 29 display autocorrelations of returns and absolute returns for tokens listed in Table 1, using all available historical data. These figures support the main text findings: return autocorrelations are mostly noise except for specific lags, while absolute return autocorrelations are large and decay slowly, indicating volatility clustering.